# InterAnimate: Taming Region-aware Diffusion Model for Realistic Human Interaction Animation

### v1

@@ Abstract
Figure 1: The highly realistic interaction animations generated from InterAnimate.
1 Introduction
Recent advances in video generation have significantly advanced human animation research, enabling increasingly realistic videos with broad applications in virtual reality, gaming, and entertainment. Popular diffusion models, such as the series of image generation  [37] and the video generation model [3] of Stable Diffusion, have shown their effectiveness in generating high-quality images and videos. Researchers are increasingly investigating human image-to-video tasks by utilizing the architecture of these models. As a pioneer in human dance generation, Disco [45] selected U-Net as its backbone and developed two ControlNets to manage the background and human pose. Following this, AnimateAnyone [14] and MagicAnimate [52] improved the fluidity of the animations by integrating motion modules into the U-Net. To fully leverage the video priors learned from large-scale video datasets, recent researches [61, 53, 12] adopted the Stable Video Diffusion as their backbone. However, most existing work focuses on isolated actions, such as dancing, and largely overlooks interactive motions—particularly those involving hands and faces. These interactions are crucial for next-generation biometric authentication systems [1, 34], which rely on dynamic, interaction-based anti-spoofing measures (e.g., “rub your chin,” “cover your left eye”) that demand high-fidelity interactive motion generation. Despite their importance, such interactions remain underexplored in the current research landscape.
To address this gap, we introduce a novel paradigm for animating realistic hand-face interactions. Unlike previous work that focuses on isolated movements or human-object interactions [50] shown in Fig. 2, our approach focuses on the complex dynamics and deformation effects that emerge from physical contact between hands and faces. Our goal is to achieve highly realistic, natural interactive animations that faithfully capture the nuances of hand-face contact.
@@ 1 Introduction
Building on InterHF dataset, we propose InterAnimate, a novel region-aware diffusion model specifically designed for interaction animation. InterAnimate introduces a region-aware interaction system that significantly enhances its ability to model the intricate dynamics of hand-face contact. Central to this system are learnable interaction latents, which include both spatial and temporal components designed to capture rich interaction priors. To seamlessly integrate these priors into the denoising process, we develop a specialized region attention block that operates in three key stages: soft quantization, cross-attention, and masking. To maintain identity consistency throughout the hand-face interactions, we implement an ID Preserver, which utilizes facial embeddings from ArcFace [7]. Additionally, we incorporate an orthogonality loss to encourage independence among the interaction latents, thereby increasing their capacity to learn a diverse array of interaction patterns. These innovations collectively enable InterAnimate to produce animations that not only achieve realistic and accurate contact, but also maintain smooth, coherent motion over time.
Figure 2: Comparison of three types of human animation video generation: human dance, focused on body movements; human-object interaction, centered on interactions between people and objects; and hand-face interaction, highlighting hand-to-face dynamics.
@@
2 Related Work
2.1 Diffusion for Video Generation
@@
In recent years, the diffusion paradigm has stimulated prosperity in numerous fields [19, 27, 43, 42, 26, 24, 36, 9, 25, 6]. Recent advances in video generation have extended diffusion models from static image synthesis to dynamic sequence modeling. Early video diffusion approaches [39] pioneered temporal modeling through 3D U-Net architectures and spatio-temporal attention mechanisms. Subsequent works like Video Diffusion Models [10] introduced cascaded super-resolution pipelines to generate high-quality videos, while Tune-A-Video [48] demonstrated one-shot video adaptation capabilities through parameter-efficient fine-tuning. The emergence of Stable Video Diffusion [4] established new benchmarks in temporal consistency through large-scale video dataset training and frame-wise latent alignment techniques. While models like HunyuanVideo [19] and SORA [27] demonstrate the potential of large-scale training for generating long-form coherent videos, current approaches remain limited in modeling fine-grained human interactions. Existing frameworks predominantly focus on full-body dance generation [20] or rigid object manipulation, neglecting the nuanced dynamics of expressive part interactions such as hand-face coordination. In this work, we bridge this gap by developing a region-aware interaction diffusion framework that combines global pose guidance with localized attention mechanisms, enabling synchronized control over both macro-body movements and micro-level facial-hand interactions while maintaining temporal coherence across extended video sequences.
2.2 Pose-guided Human Image Animation
@@
Pose-guided human image animation emphasizes the generation of realistic human movements by utilizing pose information as conditional guidance. Recent advancements in this field have been propelled by the integration of diffusion models [18, 47, 56, 17, 32, 33, 54], particularly those built on the Stable Diffusion framework [37, 3], which produce coherent animations through iterative refinement of motion dynamics. The pose condition is typically represented as a sequence of keypoints or skeletons, with features incorporated into the denoising network via ControlNet [59] or ReferenceNet [15]. Pioneering works like Disco [45] established the paradigm of combining U-Net architectures with dedicated ControlNets for background and pose management, while subsequent innovations such as AnimateAnyone [14] and MagicAnimate [52] enhanced motion fluidity through specialized motion modules. To refine the animation, some methods  [62, 44, 29] utilize the rendering maps of 3D human representation, like depth map and normal map, as the geometry-aware pose condition. Current methodologies typically address two interconnected goals: preserving identity features from reference images and synthesizing pose-driven motions. While state-of-the-art approaches [40] employ spatial-temporal attention mechanisms to separate appearance attributes from motion dynamics, recent trends [20, 61] show increasing adoption of video diffusion architectures like Stable Video Diffusion [4] to leverage large-scale video priors. However, these works predominantly focus on full-body dance generation, neglecting nuanced interactions between expressive body parts such as hands and facial features. Building upon these foundations, our method introduces a region-aware interaction system that explicitly models inter-limb coordination and hand-face synchronization through decomposed spatial-temporal attention, enabling more natural interactions in complex animation scenarios.
2.3 Human Animation with Interaction
@@
The goal of human animation with interaction is to create visual content that accurately depicts the relationships between humans and objects. While current methods primarily concentrate on human-object interaction (HOI) generation [50] or object swapping in hand-centric videos, existing approaches lack the capability to produce videos with sufficient degrees of freedom for expressive body part interactions. This limitation becomes particularly evident in hand-face interaction scenarios that require simultaneous management of facial deformation, partial occlusion, and identity preservation during contact. Recent advances in human-object interaction generation have facilitated the synthesis of rigid object manipulation [23, 35, 49], but remain inadequate for modeling soft tissue dynamics. Traditional 3D object representations [8] fail to capture non-rigid facial deformations, while spatial editing methods like HOI-Swap [55] neglect topological changes during hand-face contact. Three fundamental limitations persist across existing approaches: (i) absence of physics-inspired facial distortion models during finger contact, (ii) inability to maintain identity consistency under dynamic hand-induced occlusions, and (iii) restricted focus on isolated interactions rather than coordinated full-body motions. To address these challenges, we introduce the InterHF dataset, which contains 90,000 high-quality interaction videos. Our approach enables the generation of anchor-style videos with natural hand-face synchronization, establishing new benchmarks for complex animation with interaction.
Figure 3: The framework of InterAnimate.
3 InterHF Dataset Collection
@@
To advance research in hand-face interaction synthesis, we introduce InterHF, a large-scale, multi-modal dataset offering several distinct advantages. InterHF captures a diverse range of environmental conditions, including varied lighting setups and multiple backgrounds. The dataset also incorporates significant photometric diversity—such as specular highlights and shadows—alongside geometric variability in hand and face shapes from a balanced demographic sample. Moreover, it reflects natural behavioral patterns, including variations in interaction speed, contact force, and motion dynamics. This makes InterHF uniquely valuable for developing robust, real-world interaction synthesis models, filling a critical gap in human-centric animation research. In the following, we detail the data collection protocol, interaction taxonomy, and dataset statistics.
3.1 Data Collection Protocol
@@ 3.3 Dataset Statistics
The dataset’s extensive scale and detailed annotations, combined with its photometric and geometric diversity, offer a significant improvement over existing datasets such as TikTok [16] and Ted-talk  [38] dataset. By providing a wide range of natural behavioral variations and robust ground-truth data, InterHF sets a new benchmark for real-world interaction synthesis research.
4 Methodology
As illustrated in Fig. 3, InterAnimate is composed of two core modules that jointly enable the generation of realistic hand-face interaction animations: the Region-aware Interaction System and the ID Preserver. The Region-aware Interaction System introduces learnable interaction latents and a region attention mechanism to effectively model complex spatio-temporal contact patterns. Meanwhile, the ID Preserver ensures facial identity consistency by leveraging ArcFace embeddings and integrating them into the animation process. Together, these modules allow InterAnimate to produce animations that maintain both natural motion coherence and visual fidelity.
4.1 Region-aware Interaction System
The Region-aware Interaction System forms the backbone of InterAnimate’s animation process. It introduces two sets of two sets of Learnable Interaction Latents, including spatial latents and temporal latents, to capture diverse interaction priors. These latents learn from training samples to represent distinct hand-face interaction patterns over time. To integrate these interaction priors into the animation process, we propose the Region Attention Block to refine latent features, enrich spatial and temporal representations, and focus computation on interactive regions, enabling the video diffusion model to produce anatomically accurate and visually coherent animations.
Learnable Interaction Latents. There are lots of works that utilize learnable parameters to extract specific features [22] or to learn motion patterns  [58, 51, 28]. Inspired by this, we construct two sets of learnable vectors, dubbed Learnable Interaction Latents in our paper, to learn the interaction priors from the training samples. Specifically, spatial latents $\mathbf{L}^{\mathrm{s~{}}}\in\mathbb{R}^{n\times d}$ are used to learn spatial features such as the topology and texture of hands and faces, as well as the the appearance of the contact area. The temporal latents $\mathbf{L}^{\mathrm{t~{}}}\in\mathbb{R}^{m\times d}$ are used to learn temporal features such as interactive action patterns and contact deformation processes. The $n$ and $m$ denote the number of learnable vectors and $d$ denotes the vector dimension. During training, the learnable interaction latents can continuously learn from interaction samples and self-update, resembling a process of acquiring knowledge about interactions.
@@ 4.1 Region-aware Interaction System
0:  hidden states $\text{hidden states }\mathbf{V}^{\mathrm{i}}\in\mathbb{R}^{b\times f\times h\times w\times c}$, interaction latents $\text{interaction latents }\mathbf{L}\in\mathbb{R}^{n\times d}$, temperature $\text{temperature }\tau\in\mathbb{R}$
0:  quantized hidden states $\text{quantized hidden states }\mathbf{\bar{V}}$
1:  $\mathbf{V}_{\mathrm{flatten}}=\mathbf{V}^{\mathrm{i}}.\text{view}(-1,c))$triangleright reshape to (b*f*h*w, c)
@@
3:  $w=\text{softmax}(-D/\tau,dim=1)$
4:  $out=\text{matmul}(w,\mathbf{L})$triangleright weighted sum
5:  $\mathbf{\bar{V}}=out.\text{view}(b,f,h,w,c))$
@@
Firstly, we perform a quantization step on the input hidden states, denoted as $\mathbf{V}^{\mathrm{i}}$, which results in the generation of spatially quantized hidden states $\mathbf{\bar{V}}^{\mathrm{s}}$ and temporally quantized hidden states $\mathbf{\bar{V}}^{\mathrm{t}}$. Unlike the traditional hard indexing approach found in codebooks, this quantization method employs a soft nearest neighbor mechanism, as detailed in Algorithm 1. The proposed soft quantization approach provides several key advantages: (i) By enabling each hidden state to draw on multiple interaction latents, it enriches the representation of interaction information, thereby improving adaptability to complex patterns and enhancing generalization. (ii) The method promotes smooth transitions between states, avoiding abrupt shifts and making it easier to identify latent movement patterns. (iii) Its differentiable nature ensures a more stable and efficient training process.
@@
$\mathbf{\hat{V}}^{\mathrm{s}}$ $=\mathrm{SpatialCrossAttn}(\mathbf{\bar{V}}^{\mathrm{s}},\mathbf{L}^{\mathrm{s}},\mathbf{L}^{\mathrm{s}}),$ (1)
$\mathbf{\hat{V}}^{\mathrm{t}}$ $=\mathrm{TemporalCrossAttn}(\mathbf{\bar{V}}^{\mathrm{t}},\mathbf{L}^{\mathrm{t}},\mathbf{L}^{\mathrm{t}}),$ (2)
$\mathbf{V}^{\mathrm{latent}}=\mathrm{Mixer}(\mathbf{\hat{V}}^{\mathrm{s}},\mathbf{\hat{V}}^{\mathrm{t}})=\alpha\cdot\mathbf{\hat{V}}^{\mathrm{s}}+(1-\alpha)\cdot\mathbf{\hat{V}}^{\mathrm{t}},$ (3)
$\mathrm{SpatialCrossAttn}$ and $\mathrm{TemporalCrossAttn}$ operations all using the standard cross-attention mechanism $\mathrm{CrossAttn}$. Specifically, they follow the formula:
@@
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ denote the query, key, and value, respectively. In our implementation, the hidden states serve as the query, while the interaction latents provide both the key and the value. The operations $\mathrm{SpatialCrossAttn}$, $\mathrm{TemporalCrossAttn}$ and $\mathrm{Mixer}$ are adapted from Stable Video Diffusion  [3],
Next we utilize hand and face masks to identify the interaction-related regions, which can be formulated as
$\mathbf{V}^{\mathrm{latent}}_{\mathrm{m}}=\mathbf{V}^{\mathrm{latent}}\cdot\mathrm{M}_{\mathrm{h}}\cdot\mathrm{M}_{\mathrm{f}},$ (4)
where $\mathrm{M}_{\mathrm{h}}$ and $\mathrm{M}_{\mathrm{f}}$ are extracted from DW-poses [57] of the driving video. The application of masking is aimed at focusing on the hand and face regions while avoiding disruption to non-interactive areas, thereby accelerating the convergence of the model. Now, we have obtained the region-aware interaction states $\mathbf{V}^{\mathrm{latent}}_{\mathrm{m}}$, which captures the spatial and temporal characteristics of the interaction, guiding the denoising process. Finally, the output of the region attention block can be defined as
$\mathbf{V}^{\mathrm{o}}_{att}=\mathbf{V}^{\mathrm{i}}+\mathbf{V}^{\mathrm{latent}}_{\mathrm{m}}.$ (5)
4.2 ID Preserver
In prior research centered on image-driven generation, the majority of approaches  [3, 61, 15, 31, 30] typically utilize the CLIP image encoder for cross-attention. However, this design is inadequate for effectively maintaining identity consistency in human-centric video generation. To tackle this challenge, we propose an ID preserver to inject ID information into the hidden states. Notably, we employ face embeddings $\mathrm{A}_{\mathrm{face}}$ extracted via ArcFace  [7]. The functionality of the ID preserver can be formulated as
$\mathrm{A}_{\mathrm{emb}}=W(\mathrm{A}_{\mathrm{face}}),$ (6)
$\mathbf{V}^{\mathrm{face}}=\mathrm{CrossAttn}(\mathbf{V}^{\mathrm{i}},\mathrm{A}_{\mathrm{emb}},\mathrm{A}_{\mathrm{emb}}),$ (7)
$\mathbf{V}^{\mathrm{o}}_{\mathrm{face}}=\mathbf{V}^{\mathrm{face}}\cdot\mathrm{M}_{\mathrm{f}},$ (8)

### v2

@@ Abstract.
Human animation, video generation, interaction
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755532isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer vision
1. Introduction
Figure 1. Comparison of three types of human animation video generation: human dance, human-object interaction, and hand-face interaction.
Recent advances in video generation have significantly advanced human animation research, enabling increasingly realistic videos with broad applications in virtual reality, gaming, and entertainment. Popular diffusion models, such as the series of image generation  (Rombach et al., 2021) and the video generation model (Blattmann et al., 2023a) of Stable Diffusion, have shown their effectiveness in generating high-quality images and videos. Researchers are increasingly investigating human image-to-video tasks by utilizing the architecture of these models. As a pioneer in human dance generation, Disco (Wang et al., 2024a) selected U-Net as its backbone and developed two ControlNets to manage the background and human pose. Following this, AnimateAnyone (Hu, 2024a) and MagicAnimate (Xu et al., 2024c) improved the fluidity of the animations by integrating motion modules into the U-Net. More recently, to fully leverage the video priors learned from large-scale video datasets, MimicMotion (Zhang et al., 2024), Dispose (Li et al., 2024b), and StableAnimator (Tu et al., 2025) adopted the Stable Video Diffusion as their backbone. However, most existing work focuses on isolated actions, such as dancing, and largely overlooks interactive motions—particularly those involving hands and faces. These interactions are crucial for next-generation biometric authentication systems (Anthony et al., 2021; Muhammad et al., 2024), which rely on dynamic, interaction-based anti-spoofing measures (e.g., “rub your chin,” “cover your left eye”) that demand high-fidelity interactive motion generation. Despite their importance, such interactions remain underexplored in the current research landscape.
To address this gap, we introduce a novel paradigm for animating realistic hand-face interactions. Unlike previous work that focuses on isolated movements or human-object interactions (Xu et al., 2024a), as shown in Fig. 1, our approach focuses on the complex dynamics and deformation effects that emerge from physical contact between hands and faces. Our goal is to achieve highly realistic, natural interactive animations that faithfully capture the nuances of hand-face contact.
@@ 1. Introduction
Figure 2. The highly realistic interaction animations generated from InterAnimate.
Building on InterHF dataset, we propose InterAnimate, a novel region-aware diffusion model specifically designed for interaction animation. InterAnimate introduces a region-aware interaction system that significantly enhances its ability to model the intricate dynamics of hand-face contact. Central to this system are learnable interaction latents, which include both spatial and temporal components designed to capture rich interaction priors. To seamlessly integrate these priors into the denoising process, we develop a specialized region attention block that operates in three key stages: soft quantization, cross-attention, and masking. To maintain identity consistency throughout the hand-face interactions, we implement an ID Preserver, which utilizes facial embeddings derived from ArcFace (Deng et al., 2019). Additionally, we incorporate an orthogonality loss to encourage independence among the interaction latents, thereby increasing their capacity to learn a diverse array of interaction patterns. These innovations collectively enable InterAnimate to produce animations that not only achieve realistic and accurate contact, but also maintain smooth, coherent motion over time, as illustrated in Fig. 2.
@@
2. Related Work
2.1. Diffusion for Video Generation
@@
Early video diffusion approaches (Singer et al., 2022) pioneered temporal modeling through 3D U-Nets and spatio-temporal attention. Subsequent works like Video Diffusion Models (He et al., 2022) introduced cascaded super-resolution pipelines for quality enhancement, while Tune-A-Video (Wu et al., 2023a) enabled one-shot video adaptation through parameter-efficient fine-tuning. The emergence of Stable Video Diffusion (Blattmann et al., 2023b) established new benchmarks in temporal consistency via large-scale video dataset training and frame-wise latent alignment. While models like HunyuanVideo (Kong et al., 2024) and SORA (Lin et al., 2024a) demonstrate the potential of large-scale training for generating long-form coherent videos, current methods largely focus on full-body dance (Li et al., 2024b) or rigid object manipulation, neglecting the dynamics of expressive part interactions such as hand-face coordination. In this work, we bridge this gap by developing a region-aware interaction diffusion framework that combines global pose guidance with localized attention mechanisms, enabling synchronized control of both macro-body movements and micro-level facial-hand interactions while maintaining temporal coherence across video sequences.
2.2. Pose-guided Human Image Animation
@@
Pose-guided human animation leverages pose information to generate realistic motion, with recent advances propelled by diffusion models (Khachatryan et al., 2023; Wu et al., 2023b; Yang et al., 2023b), particularly those built on the Stable Diffusion framework (Rombach et al., 2021; Blattmann et al., 2023a), which refine motion dynamics iteratively. Poses are typically encoded as keypoints or skeletons and integrated via ControlNet (Zhang et al., 2023a) or ReferenceNet (Hu, 2024b). Pioneering works like Disco (Wang et al., 2024a) established the paradigm of combining U-Net architectures with dedicated ControlNets for background and pose management, while subsequent innovations such as AnimateAnyone (Hu, 2024a) and MagicAnimate (Xu et al., 2024c) enhanced motion fluidity through specialized motion modules. Existing methods aim to preserve identity features from reference images and synthesize pose-driven motions. While state-of-the-art approaches (Tu et al., 2025) employ spatial-temporal attention mechanisms to separate appearance attributes from motion dynamics, recent trends (Li et al., 2024b; Zhang et al., 2024) show increasing adoption of video diffusion architectures like Stable Video Diffusion (Blattmann et al., 2023b) to leverage large-scale video priors. However, these works predominantly focus on full-body dance generation, neglecting nuanced interactions between expressive body parts such as hands and facial features. Building upon these foundations, our method introduces a region-aware interaction system that explicitly models hand-face interactions in complex animation scenarios.
2.3. Human Animation with Interaction
@@
Human animation with interaction aims to depict realistic relationships between humans and objects. While current methods primarily concentrate on human-object interaction (HOI) generation (Xu et al., 2024a) or object swapping in hand-centric videos, they lack sufficient degrees of freedom for expressive body part interactions. This limitation becomes particularly evident in hand-face interaction scenarios that require simultaneous management of facial deformation, partial occlusion, and identity preservation during contact. Recent advances in HOI generation have facilitated the synthesis of rigid object manipulation (Li et al., 2023b; Peng et al., 2023; Xu et al., 2023), but remain inadequate for modeling soft tissue dynamics. Traditional 3D object representations (Ghosh et al., 2023) fail to capture non-rigid facial deformations, while spatial editing methods like HOI-Swap (Xue et al., 2024a) neglect topological changes during hand-face contact. Three fundamental limitations persist across existing approaches: (i) absence of physics-inspired facial distortion models during finger contact, (ii) inability to maintain identity consistency under dynamic hand-induced occlusions, and (iii) restricted focus on isolated interactions rather than coordinated full-body motions. To address these challenges, we introduce the InterHF dataset, which contains 90,000 high-quality interaction videos. Our approach enables the generation of natural hand-face interaction, establishing new benchmarks for complex animation with interaction.
Figure 3. The framework of InterAnimate.
3. InterHF Dataset Collection
@@
To advance research in hand-face interaction synthesis, we introduce InterHF, a large-scale, multi-modal dataset offering several distinct advantages. InterHF captures a diverse range of environmental conditions, including varied lighting setups and multiple backgrounds. The dataset also incorporates significant photometric diversity—such as specular highlights and shadows—alongside geometric variability in hand and face shapes from a balanced demographic sample. Moreover, it reflects natural behavioral patterns, including variations in interaction speed, contact force, and motion dynamics. This makes InterHF uniquely valuable for developing robust, real-world interaction synthesis models, filling a critical gap in human-centric animation research. Next, we detail the data collection protocol, interaction taxonomy, and dataset statistics.
3.1. Data Collection Protocol
@@ 3.3. Dataset Statistics
The dataset’s extensive scale and detailed annotations, combined with its photometric and geometric diversity, offer a significant improvement over existing datasets such as TikTok (Jafarian and Park, 2021) and Ted-talk  (Siarohin et al., 2021) dataset. By providing a wide range of natural behavioral variations and robust ground-truth data, InterHF sets a new benchmark for real-world interaction synthesis research.
4. Methodology
InterAnimate is composed of two core modules that jointly enable the generation of realistic hand-face interaction animations: the Region-aware Interaction System and the ID Preserver. The Region-aware Interaction System introduces learnable interaction latents and a region attention mechanism to effectively model complex spatio-temporal contact patterns. Meanwhile, the ID Preserver ensures facial identity consistency by leveraging ArcFace embeddings and integrating them into the animation process. Together, these modules allow InterAnimate to produce animations that maintain both natural motion coherence and visual fidelity.
4.1. Region-aware Interaction System
The Region-aware Interaction System forms the backbone of InterAnimate’s animation process. It introduces two sets of Learnable Interaction Latents, including spatial latents and temporal latents, to capture diverse interaction priors. These latents learn from training samples to represent distinct hand-face interaction patterns over time. To integrate these interaction priors into the animation process, we propose the Region Attention Block to refine latent features, enrich spatial and temporal representations, and focus computation on interactive regions, enabling anatomically accurate and visually coherent animation generation.
Learnable Interaction Latents. There are lots of works that utilize learnable parameters to extract specific features (Li et al., 2023a) or to learn motion patterns  (Zhang et al., 2023b; Xu et al., 2024b; Lin et al., 2024c). Inspired by this, we construct two sets of learnable vectors, dubbed Learnable Interaction Latents in our paper, to learn the interaction priors from the training samples. Specifically, spatial latents $\mathbf{L}^{\mathrm{s~}}\in\mathbb{R}^{n\times d}$ are used to learn spatial features such as the topology and texture of hands and faces, as well as the the appearance of the contact area. The temporal latents $\mathbf{L}^{\mathrm{t~}}\in\mathbb{R}^{m\times d}$ are used to learn temporal features such as interactive action patterns and contact deformation processes. The $n$ and $m$ denote the number of learnable vectors and $d$ denotes the vector dimension. During training, the learnable interaction latents can continuously learn from interaction samples and self-update, resembling a process of acquiring knowledge about interactions.
@@ 4.1. Region-aware Interaction System
[…]
@@
[…]
@@
Firstly, we perform a quantization step on the input hidden states, denoted as $\mathbf{V}^{\mathrm{i}}$, which results in the generation of spatially quantized hidden states $\mathbf{\bar{V}}^{\mathrm{s}}$ and temporally quantized hidden states $\mathbf{\bar{V}}^{\mathrm{t}}$. Unlike the traditional hard indexing approach found in codebooks, this quantization method employs a soft nearest neighbor mechanism, as detailed in Algorithm.1. The proposed soft quantization approach provides several key advantages: (i) By enabling each hidden state to draw on multiple interaction latents, it enriches the representation of interaction information, thereby improving adaptability to complex patterns and enhancing generalization. (ii) it promotes smooth transitions between states, avoiding abrupt shifts and making it easier to identify latent movement patterns. (iii) Its differentiable nature ensures a more stable and efficient training process.
@@
(1) $\mathbf{\hat{V}}^{\mathrm{s}}$ $=\mathrm{SpatialCrossAttn}(\mathbf{\bar{V}}^{\mathrm{s}},\mathbf{L}^{\mathrm{s}},\mathbf{L}^{\mathrm{s}}),$
(2) $\mathbf{\hat{V}}^{\mathrm{t}}$ $=\mathrm{TemporalCrossAttn}(\mathbf{\bar{V}}^{\mathrm{t}},\mathbf{L}^{\mathrm{t}},\mathbf{L}^{\mathrm{t}}),$
(3) $\mathbf{V}^{\mathrm{latent}}=\mathrm{Mixer}(\mathbf{\hat{V}}^{\mathrm{s}},\mathbf{\hat{V}}^{\mathrm{t}})=\alpha\cdot\mathbf{\hat{V}}^{\mathrm{s}}+(1-\alpha)\cdot\mathbf{\hat{V}}^{\mathrm{t}},$
$\mathrm{SpatialCrossAttn}$ and $\mathrm{TemporalCrossAttn}$ operations all using the standard cross-attention mechanism $\mathrm{CrossAttn}$:
@@
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ denote the query, key, and value, respectively. In our implementation, the hidden states serve as the query, while the interaction latents provide both the key and the value. The operations $\mathrm{SpatialCrossAttn}$, $\mathrm{TemporalCrossAttn}$ and $\mathrm{Mixer}$ are adapted from Stable Video Diffusion  (Blattmann et al., 2023a),
To further refine $\mathbf{V}^{\mathrm{latent}}$, we utilize hand and face masks to identify the interaction-related regions, which can be formulated as
(4) $\mathbf{V}^{\mathrm{latent}}_{\mathrm{m}}=\mathbf{V}^{\mathrm{latent}}\cdot\mathrm{M}_{\mathrm{h}}\cdot\mathrm{M}_{\mathrm{f}},$
where $\mathrm{M}_{\mathrm{h}}$ and $\mathrm{M}_{\mathrm{f}}$ are extracted from DW-poses (Yang et al., 2023a) of the driving video. The application of masking is aimed at focusing on the hand and face regions while avoiding disruption to non-interactive areas, thereby accelerating the convergence of the model. Now, we have obtained the region-aware interaction states $\mathbf{V}^{\mathrm{latent}}_{\mathrm{m}}$, which captures the spatial and temporal characteristics of the interaction, guiding the denoising process. Finally, the output of the region attention block can be defined as
(5) $\mathbf{V}^{\mathrm{o}}_{att}=\mathbf{V}^{\mathrm{i}}+\mathbf{V}^{\mathrm{latent}}_{\mathrm{m}}.$
4.2. ID Preserver
In prior research centered on image-driven generation, the majority of approaches  (Blattmann et al., 2023a; Zhang et al., 2024; Hu, 2024b) typically utilize the CLIP image encoder for cross-attention. However, this design is inadequate for effectively maintaining identity consistency in human-centric video generation. To tackle this challenge, we propose an ID preserver to inject ID information into the hidden states. Notably, we employ face embeddings $\mathrm{A}_{\mathrm{face}}$ extracted via ArcFace  (Deng et al., 2019). The functionality of the ID preserver can be formulated as
(6) $\mathrm{A}_{\mathrm{emb}}=W(\mathrm{A}_{\mathrm{face}}),$
(7) $\mathbf{V}^{\mathrm{face}}=\mathrm{CrossAttn}(\mathbf{V}^{\mathrm{i}},\mathrm{A}_{\mathrm{emb}},\mathrm{A}_{\mathrm{emb}}),$
(8) $\mathbf{V}^{\mathrm{o}}_{\mathrm{face}}=\mathbf{V}^{\mathrm{face}}\cdot\mathrm{M}_{\mathrm{f}},$